a photogrpaher wishes to mix 2 liters of a 5 percent acetic acid solution with a 10 percent solution to get a 7 percent solution. How many liters of 10 percent solution must be added?
Answer by richwmiller(17219)

0.1*x+0.05*2=0.07(2+x)
0.1*x+0.1=0.14+0.07x
0.1x-0.07x=0.14-0.1
0.03x=0.04
x=1.33333333 liter of 10% is added to 2 liter of 5% to get 3.33333333 liter at 7%
check
0.1*1.33333333+0.05*2=0.07(2+1.33333333)
0.13333333+0.1=0.07(3.33333333)
0.23333333=0.23333333
ok